OpenStudy (anonymous):

How does completing the trinomial square help you solve an equation by using the principle of square root? Can you consider the following and demonstrate the concept? x^2 + 4x = 21

OpenStudy (anonymous):

\[x^2+4x=21\] \[x^2+4x+4=25\] \[(x+4)^2=25\] \[x+4=4\] or \[x+4=-5\]

OpenStudy (anonymous):

the picture goes like this. draw a square with side x. its area is \[x^2\] attach to two sides a rectangle with side 2. each has area \[2x\] for a total of \[x^2+4x\] this picture is not a square. to make it a square you have to actually complete the square, but adding a little \[2\times 2\] square for the piece that is missing.

OpenStudy (anonymous):

since that little square has area 4 you now have a total of 21+4 = 25 and you have a ' perfect square" whose sides are now \[x+2\]

OpenStudy (anonymous):

oops guess that didn't work.

OpenStudy (anonymous):

lol....I thought what the heck is going on?